# Strategic Pricing and Ranking in Recommendation Systems with Seller Competition

### v1

@@ 7.1 Equal sellers: Nash or $\epsilon$-Nash equilibria
We now consider ‘near symmetric’ environments, by considering the cases where the sellers are relatively similar in their reputation or strength, represented by almost equal $\{\eta_{a}\}_{a}$ values in (22). This game admits the typical solution in the form of a NE or an $\epsilon$-NE. The following theorem formally characterizes two types of scenarios and identifies the corresponding equilibria (proof is provided in C).
@@
    Further, assume that $(\eta_{1}-\eta_{M+1})<\epsilon\min\left\{e\min_{k\leq M}\alpha_{k},1\right\}.$ Then the strategy profile
    ${\bm{\delta}}_{*}=(\eta_{2},\eta_{3},\cdots,\eta_{M+1},\eta_{M+1},\delta_{M+2},\cdots,\delta_{N})$
    is an $\epsilon$-equilibrium, if $\delta_{k}\leq\eta_{k}$ for all $k>(M+1)$.
@@
When we attempt to adapt the framework of [27] for analysis, we encountered a problem on a set of measure zero; it is amply clear in the literature that sets of measure zero often do not alter the outcomes or the consequences. In this paper, we manage to extend the notion of EC to navigate through such measure-zero violations and establish that our unified seller game posses the modified EC as a solution. We now take a brief detour to explain the framework of [27], and then to extend the notions.
@@ 7.3 Unequal sellers equilibrium—equilibrium cycle
Define the product measure $\mu=\prod_{k}\mu_{k}$ with $\mu_{k}=\mu_{leb}$, the standard Lebesgue measure for all $k\leq M$ and $\mu_{k}={\Delta({\eta_{k}})}$, the Dirac measure concentrated at $\eta_{k}$ for each $k>M$. We now claim and prove that the set defined in (27) is the $\mu$-EC (proof is in C).
@@
From the platform’s perspective, this outcome is less favourable compared to the outcome with equal or near-equal sellers (see Theorem 2).In the latter case the sellers operate at (or near) their marginal costs and the platform extracts the maximum possible commission from each of them. While with unequal sellers (especially when $\eta_{M+1}$ is significantly smaller) there are two troubles: i) the sellers do not settle or a fixed price, the $\mu$-EC solution suggests the unrest with EC components; and ii) the sellers may often quote commissions that are near the threat point $\eta_{M+1}$, which can imply a significant reduction in the revenue of the platform. Basically a bigger seller competition with more competitive sellers imply a bigger advantage for the platform.
@@ 7.4 Numerical Results
Figure 3: BR dynamics of seller commissions in the unified seller game with $N=5$, $M=3$, and $\gamma=0.1$. Horizontal dashed lines represent the theoretical thresholds (26).
Figure 3 shows the evolution of sellers’ commission strategies over iterations. The horizontal dashed lines correspond to the analytically computed threshold values ${\tilde{\eta}}_{k}$ (defined in (26)) for the top 3 positions in the menu. As seen in the figure, the top three sellers might not secure positions in the menu in the initial iteration, but soon reach a position where they persistently ensure inclusion in the top-$3$ menu. Further, they do not stabilize to fixed commission levels, but instead oscillate within the interval $\eta_{4},{\tilde{\eta}}_{k}]$. As seen these strategic interactions have a pattern: initially, all three sellers undercut one another until seller 3 hits the threshold ${\tilde{\eta}}_{3}$; subsequently, sellers 1 and 2 continue undercutting each other until both reach ${\tilde{\eta}}_{1}={\tilde{\eta}}_{2}$. Once all three have reached their respective thresholds, the dynamics restarts from $\eta_{4}=0.6$, by one of them making the first move; and the cycle continues. This oscillatory pattern reflects the structure of the equilibrium cycle shown in Theorem 3. In contrast, after some iterations (once the top $3$ sellers reach above $\eta_{4}$), sellers 4 and 5, persistently fail to grab a position in the menu. This aligns with the Theorem 3, which states that at equilibrium, NMN-M sellers do not get position in the $\mu$-connected equilibrium cycle.
@@ 8 Conclusions
Interestingly, we show that the standard pure-strategy Nash equilibria (NE) exists only for the scenarios with equal or near equal sellers; more interestingly at the resulting NE the sellers operate at marginal costs, exactly like in Bertrand duopoly. Once the sellers are of different strengths, the outcome of the game in the space of pure strategies, is captured by a novel solution concept, namely $\mu$-connected equilibrium cycle. Our analysis also reveals how weaker sellers can act as strategic threats to stronger ones, pushing the latter to offer higher commissions in order to maintain visibility.
@@ 1. $\bullet$
    Now player $1$ can appropriately choose $\delta_{1}^{\prime}\in(\delta_{k},{\tilde{\eta}}_{2})$ to improve its utility as:
    ${U}_{1}(\delta_{1},{\bm{\delta}}_{-1})\stackrel{{\scriptstyle a}}{{\leq}}\frac{\left(\eta_{1}-\delta_{1}\right)}{e\alpha_{1}}\frac{\sum_{0\leq j<{b}}{\tilde{\gamma}}^{j}\mathds{1}_{\left\{j\leq M\right\}}}{b}<\frac{\left(\eta_{1}-\delta_{1}^{\prime}\right)}{e\alpha_{1}}={U}_{1}(\delta_{1}^{\prime},{\bm{\delta}}_{-1}).$
    In the above, if $\delta_{1}=\delta_{k}$, we have equality in ‘$a$’; otherwise, we have a strict inequality in ‘$a$’, because for any $n>b$, (which represents the position of seller 1 in the menu with $\bm{\delta}$}),
    $\tilde{\gamma}^{n}\leq\frac{1+\tilde{\gamma}+\tilde{\gamma}^{2}+\cdots+\tilde{\gamma}^{b}}{b}<1,\ \mbox{ as }b\geq 2.$
    Also, observe ${U}_{1}(\delta_{1}^{\prime},{\bm{\delta}}_{-1})>{U}_{1}({\tilde{\eta}}_{2},{\bm{\delta}}_{-1})$ and further using (34), the remaining part of the second condition (25) is satisfied for this sub-case.
@@ 3. $\bullet$
    If $\delta_{k}>\delta_{l}$ then player $k$ can choose any $\delta_{k}^{\prime}\in(\delta_{l},\delta_{k})$ and strictly improve:
    ${U}_{k}(\delta_{k},{\bm{\delta}}_{-k})=\frac{\left(\eta_{k}-\delta_{k}\right)}{e\alpha_{k}}<\frac{\left(\eta_{k}-\delta_{k}^{\prime}\right)}{e\alpha_{k}}={U}_{k}(\delta_{k}^{\prime},{\bm{\delta}}_{-k}).$
    The second part of (25) follows as in the previous sub-case.

### v2

@@ 7.1 Equal sellers: Nash or $\epsilon$-Nash equilibria
We now consider ‘near symmetric’ environments, by considering the cases where the sellers are relatively similar in their reputation or strength, represented by almost equal $\{\eta_{a}\}_{a}$ values in (22). This game admits the typical solution in the form of a NE or an $\epsilon$-NE. The following theorem formally characterizes two types of scenarios and identifies the corresponding equilibria (proof is provided in D).
@@
    Further, assume that $(\eta_{1}-\eta_{M+1})<\epsilon\min\left\{e\min_{k\leq M}\alpha_{k},1\right\}.$ Then the strategy profile δ_* = (η_2, η_3, , η_M+1, η_M+1, δ_M+2, , δ_N) is an $\epsilon$-equilibrium, if $\delta_{k}\leq\eta_{k}$ for all $k>(M+1)$.
@@
When we attempt to adapt the framework of [6] for analysis, we encounter a problem on a set of measure zero; it is amply clear in the literature that sets of measure zero often do not alter the outcomes or the consequences. In this paper, we manage to extend the notion of EC to navigate through such measure-zero violations and establish that our unified seller game posses the modified EC as a solution. We now take a brief detour to explain the framework of [6], and then to extend the notions.
@@ 7.3 Unequal sellers equilibrium—equilibrium cycle
Define the product measure $\mu=\prod_{k}\mu_{k}$ with $\mu_{k}=\mu_{leb}$, the standard Lebesgue measure for all $k\leq M$ and $\mu_{k}={\Delta({\eta_{k}})}$, the Dirac measure concentrated at $\eta_{k}$ for each $k>M$. We now claim and prove that the set defined in (27) is the $\mu$-EC (proof is in D).
@@
From the platform’s perspective, this outcome is less favorable compared to the outcome with equal or near-equal sellers (see Theorem 2).In the latter case the sellers operate at (or near) their marginal costs and the platform extracts the maximum possible commission from each of them. While with unequal sellers (especially when $\eta_{M+1}$ is significantly smaller) there are two troubles: i) the sellers do not settle on a fixed price, the $\mu$-EC solution suggests the unrest with EC components; and ii) the sellers may often quote commissions that are near the threat point $\eta_{M+1}$, which can imply a significant reduction in the revenue of the platform. Basically a bigger seller competition with more competitive sellers imply a bigger advantage for the platform.
@@ 7.4 Numerical Results
(a) $N=5$, $M=3$
(b) $N=20$, $M=10$ (plots for sellers 12–20 omitted for clarity)
Figure 3: $\epsilon$-Best response (BR) dynamics of seller commissions in the unified seller game with $\gamma=0.1$. Horizontal dashed lines represent the theoretical thresholds (26).
[Uncaptioned image] Figure 4: The optimal menu and the price vector of a random problem with $N=18$, $M=5$ for different $\gamma$.    [Uncaptioned image] Figure 5: OS-OP game with $\gamma=0.7$. [Uncaptioned image] Figure 6: OS-OP game with $\gamma=0.1$. Comparison of BR dynamics OS-OP games with $\alpha=[0.1,0.21,0.3,0.22,0.24]$, $N=5$, $M=3$, and $c_{ma}=[1.5,0.95,0.97,0.81,1.0]$.
Figure 3 shows the evolution of sellers’ commission strategies over iterations. The horizontal dashed lines correspond to the analytically computed threshold values ${\tilde{\eta}}_{k}$ (defined in (26)) for the top $M$ positions in the menu. As seen in the Figure 3a, the top three sellers might not secure positions in the menu in the initial iteration, but soon reach a position where they persistently ensure inclusion in the top-$3$ menu. Further, they do not stabilize to fixed commission levels, but instead oscillate within the interval $\eta_{4},{\tilde{\eta}}_{k}]$. As seen these strategic interactions have a pattern: initially, all three sellers undercut one another until seller 3 hits the threshold ${\tilde{\eta}}_{3}$; subsequently, sellers 1 and 2 continue undercutting each other until both reach ${\tilde{\eta}}_{1}={\tilde{\eta}}_{2}$. Once all three have reached their respective thresholds, the dynamics restarts from $\eta_{4}=0.6$, by one of them making the first move; and the cycle continues. This oscillatory pattern reflects the structure of the equilibrium cycle shown in Theorem 3. In contrast, after some iterations (once the top $3$ sellers reach above $\eta_{4}$), sellers 4 and 5, persistently fail to grab a position in the menu. This aligns with the Theorem 3, which states that at equilibrium, NMN-M sellers do not get position in the $\mu$-connected equilibrium cycle. An exactly similar pattern appears in Figure 3b, even when the menu size (M10M-10) and the number of items ($N=20$) are large.
In Figures 5 and 6, we present yet another numerical study, where the $\epsilon$-BRs (and the platform policies) are computed without using small $\gamma$-approximation (18); basically we solve the platform optimization problem directly using the DP equations (9), while computing the BRs of sellers (in each iteration); we then compare the OP game BR-dynamics with that of OS game. The parameters of the example are provided in the caption itself. Interestingly, we again have equilibrium cycles; further the OP-BR-dynamics (almost) match with that corresponding to the OS game when $\gamma$ is small (in the Figure 6 with $\gamma=0.1$, the OP-circle-curves almost coincide with OS straight-line-curves); more interestingly, the cyclic-structure of the two dynamics are not significantly differently even for larger $\gamma$ (in the Figure 5 with $\gamma=0.7$, we again have cycles for both the cases and further the cycle lengths are almost the same).
In the final case study, we analyze the optimal policy of the platform as a function of $\gamma$. We consider several case studies with different $(N,M)$ where different instances of $\alpha,c_{ma}$ parameters are generated randomly (solution for one instance is provided in Figure 4). In all the case-studies, without exception, we make the following striking observation — the optimal menu in the OP game (computed by solving DP equations) does not depend upon $\gamma$ and further matches exactly with the one in the OS game (given by h-index rule, (19)). However, the optimal price-vectors vary with $\gamma$ (see Figure 4). Once the optimal menu is computed using the simplistic $h$-index rule, it is significantly easier to obtain the price-components of the optimal policy, while solving the DP equations (9) in backward fashion.
@@ 8 Conclusions
The inclusion of strategic influence of sellers into the traditional cascade-click models has provided several new insights in both the games. Interestingly, we show that the standard pure-strategy Nash equilibria (NE) exists only for the scenarios with equal or near equal sellers; more interestingly at the resulting NE the sellers operate at marginal costs, exactly like in Bertrand duopoly. Once the sellers are of different strengths, the outcome of the game in the space of pure strategies, is captured by a novel solution concept, namely $\mu$-connected equilibrium cycle. Our analysis also reveals how weaker sellers can act as strategic threats to stronger ones, pushing the latter to offer higher commissions in order to maintain visibility.
@@ Appendix C Proofs related to Section 6
Proof of Lemma 4.
Appendix D Proofs related to Section 7
@@ 1. $\bullet$
    Now player $1$ can appropriately choose $\delta_{1}^{\prime}\in(\delta_{k},{\tilde{\eta}}_{2})$ to improve its utility as: U_1(δ_1, δ_-1) ≤a ( η1- δ1) e α10 ≤j ¡ b~γj1{ j ≤M }b ¡ ( η1- δ1’ ) e α1 = U_1(δ_1’, δ_-1). In the above, if $\delta_{1}=\delta_{k}$, we have equality in ‘$a$’; otherwise, we have a strict inequality in ‘$a$’, because for any $n>b$, (which represents the position of seller 1 in the menu with $\bm{\delta}$}), ~γ^n ≤1 + ~γ+ ~γ2+ ⋯+ ~γbb ¡ 1, as b ≥2. Also, observe ${U}_{1}(\delta_{1}^{\prime},{\bm{\delta}}_{-1})>{U}_{1}({\tilde{\eta}}_{2},{\bm{\delta}}_{-1})$ and further using (34), the remaining part of the second condition (25) is satisfied for this sub-case.
@@ 3. $\bullet$
    If $\delta_{k}>\delta_{l}$ then player $k$ can choose any $\delta_{k}^{\prime}\in(\delta_{l},\delta_{k})$ and strictly improve: U_k(δ_k, δ_-k) = ( ηk- δk) e αk ¡ ( ηk- δk’ ) e αk = U_k (δ_k’, δ_-k). The second part of (25) follows as in the previous sub-case.